# Questioning the Survey Responses of Large Language Models

### v3

@@ Abstract
As large language models increase in capability, researchers have started to conduct surveys of all kinds on these models in order to investigate the population represented by their responses. In this work, we critically examine language models’ survey responses on the basis of the well-established American Community Survey by the U.S. Census Bureau and investigate whether they elicit a faithful representations of any human population. Using a de-facto standard multiple-choice prompting technique and evaluating 39 different language models using systematic experiments, we establish two dominant patterns: First, models’ responses are governed by ordering and labeling biases, leading to variations across models that do not persist after adjusting for systematic biases. Second, models’ responses do not contain the entropy variations and statistical signals typically found in human populations. As a result, a binary classifier can almost perfectly differentiate model-generated data from the responses of the U.S. census. At the same time, models’ relative alignment with different demographic subgroups can be predicted from the subgroups’ entropy, irrespective of the model’s training data or training strategy. Taken together, our findings suggest caution in treating models’ survey responses as equivalent to those of human populations.
@@ 1 Introduction
Many established survey questionnaires together with the carefully collected answer statistics are publicly available. Machine learning researchers have identified the potential benefits of building on this valuable data resource to study large language models (LLMs). Survey questions offer a way to systematically prompt LLMs, and the aggregate statistics over answers collected by surveying human populations serve as a reference point for evaluation. As a result, the use of surveys has recently gained popularity for studying LLMs’ biases [Santurkar et al.,, 2023; Durmus et al.,, 2023]. At the same time, researchers in the social sciences have explored using LLMs to emulate the survey responses of specific human populations [Argyle et al.,, 2022; Lee et al.,, 2023]. If effective proxies, simulated responses could augment or replace the expensive data collection process of human responses.
It is tempting to prompt LLMs with survey questions, due to their syntactic similarity to question answering tasks [Brown et al.,, 2020; Liang et al.,, 2022]. However, it is a priori unclear how to interpret their answers. Rather than knowledge testing, surveys seek to elicit aggregate statistics over individuals, providing an unbiased view on the properties of the population they represent. The quality of survey data hinges on the validity and robustness of the conclusions that can be drawn from it. Clearly, running a survey on LLMs is different from interrogating humans and comes with distinct challenges. While much research has gone into carefully designing surveys to ensure faithful human responses, it is unclear whether prompting LLMs with the same surveys satisfies similar premises out-of-the-box. We devote this work to gain insights into LLMs’ survey responses, what we can expect to learn from them, and to what extent they resemble those of human populations.
@@ 1.1 Our work
The basis of our investigation is the American Community Survey111https://www.census.gov/programs-surveys/acs (ACS), a demographic survey conducted by the U.S. Census Bureau at a national level, on a yearly basis. We curate a questionnaire by selecting 25 multiple choice questions from the 2019 ACS. We prompt 39 language models of varying size with these questions, individually and in sequence, and we record their probability distribution over answers. Based on the collected data, we investigate:
1.
What can we learn about LLMs, and their relative differences, from their survey responses?
2.
Does the data generated by prompting models to answer the ACS questionnaire resemble the census data collected by surveying the U.S. population?
@@
We start by inspecting models’ distributions over answers to individual survey questions when the questions are asked independently. We observe that the entropy of models’ responses differs substantially across models of varying size. We find that this differences arise because strong ordering and labeling biases confound models’ answers. In fact, after adjusting for such systematic biases through randomized choice ordering, we find that response distributions are very similar across models and tend to correspond to highly balanced answers.
Comparing models’ responses to those of the U.S. census population, we find that natural variations in entropy across questions are not reflected in models’ responses. Instead, on average across questions, models’ responses are no closer to the census population, or the population of any state within the US, than to a fixed uniform baseline. As a result, the relative alignment of model responses with different demographic subgroups can be explained by the entropy of the subgroups’ responses, irrespective of the data or training procedure employed to train the model. We verify that our findings generalize beyond the ACS to other surveys considered by prior work, thus providing important context to prior studies that employ surveys to examine the biases of LLMs.
Lastly, we prompt LLMs to repeatedly answer entire survey questionnaires, generating for each model a synthetic tabular dataset which emulates in form the ACS census data. We present questions in a sequential manner, keeping a model’s previous answers in context when prompting it to answer subsequent questions of the questionnaire. We then investigate whether these synthetic datasets resemble the responses collected by the U.S. Census. We use the discriminator test as an investigative tool: we construct a binary prediction task aiming to discriminate between model-generated data and the U.S. census data. We show that this prediction task can be solved with very high accuracy () for all models up to 70B parameters. For the latest GPT-4 model the accuracy is , indicating that there is a substantial gap between model-generated data and human responses.
@@
More broadly, our findings suggest caution when treating language models’ survey responses as a faithful representation of any human population, putting the burden of proof on researchers wishing to draw valid conclusions about human populations from data collected by surveying LLMs.
@@ 1.2 Related work
Despite the syntactical similarities, evaluating LLMs on the basis of their survey responses differs from traditional question answering evaluations [Liang et al.,, 2022]. Question answering (QA) tasks predominantly serve the purpose of knowledge testing [e.g., Kwiatkowski et al.,, 2019; Rajpurkar et al.,, 2016; Talmor et al.,, 2019; Mihaylov et al.,, 2018]. In such setting, a language model’s answer to some unambiguous input question is extracted by computing its most likely completion. Alternatively, models’ most likely response to questions that lack a clear answer (e.g., “Angela and Patrick are sitting together. Who is an entrepreneur?”) have been used to investigate various biases of LLMs [Li et al.,, 2020; Mao et al.,, 2021; Perez et al.,, 2022; Abid et al.,, 2021; Jiang et al.,, 2022].
@@
When evaluating LLMs on the basis of survey questions, it is not models’ most likely completion that is studied, but rather models’ probability distribution over various answer choices. Santurkar et al., [2023] study LLMs’ answer distributions for multiple-choice opinion polling questions, measuring their similarity to those of various U.S. demographic groups. They extract models’ answer distributions from the next token probabilities corresponding to each answer choice. Durmus et al., [2023] employ a similar methodology but instead consider transnational opinion surveys. We adopt this popular methodology to investigate the properties of models’ answer distributions on the basis of a well-established demographic survey, beyond measuring the relative similarity of models’ responses to the survey responses of different human populations.
@@
In addition to asking questions individually, we also prompt models to complete entire survey questionnaires. We present questions in a sequential manner, keeping a model’s previous answers in context when prompting the model to answer subsequent questions. This methodology resembles prior work by Hartmann et al., [2023]; Rutinowski et al., [2023]; Motoki et al., [2023]; Feng et al., [2023] who sequentially prompt language models to answer entire political compass or voting advice questionnaires. But instead of aggregating answers into a political affinity score, our focus is on examining whether models’ responses resemble those of human populations.
@@
Lastly, there is an emerging body of research that integrates LLMs into computational social science [Ziems et al.,, 2023]. This includes tasks such as taxonomic labeling, where language models are employed for tasks such as opinion prediction [Kim and Lee,, 2023; Mellon et al.,, 2022], and free-form coding, where language models are used to generate explanations for social science constructs [Nelson et al.,, 2021]. Recent studies have also investigated the feasibility of using LLMs to simulate human participants in psychological, psycholinguistic, and social psychology experiments [Dillion et al.,, 2023; Aher et al.,, 2023], or as proxies for specific human populations in social science research [Argyle et al.,, 2022; Lee et al.,, 2023; Sanders et al.,, 2023] and economics [Brand et al.,, 2023; Horton,, 2023]. Within this context, our work suggests caution in relying on the survey responses of LLMs to elicit synthetic responses that resemble those of human populations.
@@ 2 Surveying language models
We employ the de-facto standard methodology to survey language models introduced by Santurkar et al., [2023]. For every survey question, we collect language models’ probability distribution over answer choices. Formally, for a given model and survey question we define the model’s survey response as a categorical random variable which can take on values corresponding to the number of answer choices to question . We determine the event probabilities of by prompting model as follows:
@@
We query language models with the input prompt and obtain their output distribution over next-token probabilities. We select the output probabilities corresponding to each answer choice (e.g., the tokens “A”, “B”, etc.), and we renormalize to obtain the probability distribution over survey answers. For OpenAI’s models, we only have access to the top-5 next-token log probabilities through the OpenAI API. In this case, we assign to the unseen probabilities (if any) the minimum between the remaining probability mass and the smallest observed probability, following the methodology of Santurkar et al., [2023].
The chosen style of prompt is standard for question answering tasks [Hendrycks et al.,, 2021], used in OpinionQA [Santurkar et al.,, 2023], and follows the best practices for social science research recommended by Ziems et al., [2023]. For completeness we perform several prompt ablations, including the prompt variations used by Argyle et al., [2022], Santurkar et al., [2023] and Durmus et al., [2023]. We find our take-aways to be robust to such changes, see Appendix D.
@@
We use a representative subset of 25 multiple-choice questions from the 2019 ACS questionnaire. We denote the set of questions by . The questions cover basic demographic information, education attainment, healthcare coverage, disability status, family status, veteran status, employment status, and income. We generally consider the questions and answers as they appear in the ACS questionnaire. We refer to Appendix A.1 for the exact framing we used for each question.
@@
We survey 39 language models of size varying from 110M to 175B parameters: the base models GPT-2 [Radford et al.,, 2019], GPT-Neo [Black et al.,, 2021], Pythia [Biderman et al.,, 2023], MPT [MosaicML,, 2023], LLaMA [Touvron et al., 2023a, ], Llama 2 [Touvron et al., 2023b, ], and GPT-3 [Brown et al.,, 2020]; as well as the instruct variants of MPT 7B and GPT NeoX 20B, the Dolly fine-tune of Pythia 12B [Databricks,, 2023], the Vicuna and Koala fine-tunes of LLaMA 7B and 13B [Geng et al.,, 2023; Chiang et al.,, 2023], Llama 2 Chat [Touvron et al., 2023b, ], GPT-3.5, GPT-4 [OpenAI,, 2023], and the text-davinci variants of GPT-3 [Ouyang et al.,, 2022].
@@
We use the responses collected by the U.S. Census Bureau when surveying the U.S. population as our reference data. In particular, we use the 2019 ACS public use microdata sample222https://www.census.gov/programs-surveys/acs/microdata (henceforth census data). The data contains the anonymized responses of around 3.2 million individuals in the United States. For each survey question , we denote the census’ population-level response as a categorical random variable whose event probabilities are the relative frequency of each answer choice among survey respondents. We use to denote the uniform distribution over answers. Given these two reference points, we evaluate language models’ responses along two dimensions:
@@
For several investigations we survey models under randomized choice ordering. For a given question , we prompt models with different permutations of the answer choice ordering, i.e., the assignment of answers (e.g., “male”, “female”) to choice labels (“A”, “B”, etc), while the order of choice labels is kept alphabetically. We evaluate models’ survey responses under all possible choice orderings and we use to denote the expected distribution over answers and to denote the expected distribution over selected choice labels. For questions with more than 6 answers we evaluate a maximum of 5000 permutations. For OpenAI’s models we evaluate up to 50 permutations due to the costs of querying the OpenAI API. This distinction serves to decouple a model’s tendency towards picking a particular answer from its tendency towards picking a particular choice label. In the following we refer to the expected survey response under uniformly distributed choice ordering as the adjusted survey response. See Figure 1 for an illustration of the methodology used for adjusting for choice ordering response biases.
3 Inspecting models’ survey responses
@@
For a first investigation, we consider the normalized entropy of models’ responses to the “SEX", “HICOV”, and “FER" questions. The SEX question inquiries about the person’s sex, encoded as male female, the HICOV question inquiries whether the person is currently covered by any health insurance plan, and the FER question inquires whether the person has given birth in the past 12 months. When surveying the U.S. population, these three questions elicit responses with very different entropy; responses to the SEX question are almost uniformly distributed, whereas most people answer “No” to the FER question. In contrast, as shown in Figure 2, the entropy of models’ responses to these three questions are surprisingly similar. In particular, we find that the entropy of models’ responses tends to increase log-linearly with model size. This trend is consistent across all ACS survey questions, see Figure 9 in Appendix B.1.
For a broader picture, we illustrate models’ response entropy across all survey questions in Figure 3(a). The blue dots represent models’ responses to individual questions, and the green dots represent the entropy of the responses of the U.S. census. We order models by size. We observe that the entropy of responses of the U.S. census greatly varies across questions. In contrast, for any given model, the entropy of its responses varies substantially less so.
Overall, we find that models’ response distributions seem to be widely independent of the survey question asked, and variations across models are much larger than variations across questions. This lead us to suspect that variations across models might arise mostly due to systematic biases.

### v4

@@ Abstract
Surveys have recently gained popularity as a tool to study large language models. By comparing survey responses of models to those of human reference populations, researchers aim to infer the demographics, political opinions, or values best represented by current language models. In this work, we critically examine this methodology on the basis of the well-established American Community Survey by the U.S. Census Bureau. Evaluating 43 different language models using de-facto standard prompting methodologies, we establish two dominant patterns. First, models’ responses are governed by ordering and labeling biases, for example, towards survey responses labeled with the letter ‘A’. Second, when adjusting for these systematic biases through randomized answer ordering, models across the board trend towards uniformly random survey responses, irrespective of model size or pre-training data. As a result, in contrast to conjectures from prior work, survey-derived alignment measures often permit a simple explanation: models consistently appear to better represent subgroups whose aggregate statistics are closest to uniform for any survey under consideration.
@@ 1 Introduction
Many established survey questionnaires together with the carefully collected answer statistics are publicly available. Machine learning researchers have identified the potential benefits of building on this valuable data resource to study large language models (LLMs). Survey questions offer a way to systematically prompt LLMs, and the aggregate statistics over answers collected by surveying human populations serve as a reference point for evaluation. As a result, the use of surveys has recently gained popularity for studying LLMs’ biases [Santurkar et al.,, 2023; Durmus et al.,, 2023]. Also prompting LLMs with survey questions, researchers in the social sciences have explored using LLMs to emulate the survey responses of human populations [Argyle et al.,, 2023; Lee et al.,, 2023]. If effective proxies, simulated responses could augment or replace the expensive data collection process involving human subjects and provide insights into subpopulations that are otherwise hard to reach.
It is tempting to prompt LLMs with survey questions, due to their syntactic similarity to question answering tasks [Brown et al.,, 2020; Liang et al.,, 2022]. However, it is a priori unclear how to interpret their answers. Rather than knowledge testing, surveys seek to elicit aggregate statistics over individuals, providing an unbiased view on the properties of the population they are targeting. The quality of survey data hinges on the validity and robustness of the conclusions that can be drawn from it. Clearly, running a survey on LLMs is different from interrogating humans and thus it comes with distinct challenges. While much research has gone into carefully designing surveys to ensure faithful human responses, it is unclear whether prompting LLMs with the same surveys satisfies similar premises. We devote this work to gain systematic insights into the survey responses of LLMs, what we can expect to learn from them, and to what extent they resemble those of human populations.
@@ 1.1 Our work
The basis of our investigation is the American Community Survey111https://www.census.gov/programs-surveys/acs (ACS), a demographic survey conducted by the U.S. Census Bureau at a national level, on a yearly basis. We curate a questionnaire composing of 25 multiple choice questions from the 2019 ACS. We prompt 43 language models of varying size with these questions, individually and in sequence, and we record their probability distribution over answers. Based on the collected data, we investigate the following two questions: What can we infer about LLMs, and the data they have been trained on, from their survey responses? Does the data generated by prompting models to answer the ACS questionnaire qualitatively resemble the census data collected by surveying the U.S. population? See Figure 1.
@@
We start by inspecting models’ distributions over answers to individual survey questions when the questions are asked independently. We observe that the entropy of response distributions differs substantially across models of varying size. Entropy tends to increase log linearly with model size, and it is preserved across different questions asked. We find that this differences arise because strong ordering and labeling biases confound models’ answers. In fact, after adjusting for such systematic biases through randomized choice ordering, we find that response distributions are very similar across models and tend to correspond to highly balanced answers.
Comparing models’ adjusted responses to those of the U.S. census population, we find that natural variations in entropy across questions are not reflected in the responses. Instead, on average across questions, models’ responses are no closer to the census population, or the population of any state within the US, than to a fixed uniform baseline. This qualitative difference between model responses and human data puts into question the insights that can be gained from such comparisons. We find that even after instruction-tuning this trend persists, and model responses have consistently higher entropy than any human population we compare to, independent of the survey used. Only for models of size larger than billion parameters we can recognize a trend that the divergence between model responses and the census data decreases after instruction-tuning.
With these insights in mind, we inspect conjectures from prior work related to survey derived alignment metrics, that is, that differences in similarity between models’ and populations’ responses might be attributable to certain demographics being better represented in the training data. Instead, our results suggest a much simpler explanation: the relative alignment of model responses with different demographic subgroups can be explained by the entropy of the subgroups’ responses, irrespective of the data or training procedure employed to train the model. We demonstrate this beyond the ACS on other surveys considered by prior work. As such, our findings provide important context to prior studies that employ surveys to examine the biases of LLMs.
@@
More broadly, our findings suggest caution when treating language models’ survey responses as a faithful representation of any human population, at least a present time, as it could lead to potentially misguided conclusions about alignment.
@@ 1.2 Related work
Despite the syntactical similarities, there are important differences between evaluating LLMs on the basis of their survey responses and traditional question answering evaluations [Liang et al.,, 2022]. Question answering (QA) tasks predominantly serve the purpose of knowledge testing [e.g., Kwiatkowski et al.,, 2019; Rajpurkar et al.,, 2016; Talmor et al.,, 2019; Mihaylov et al.,, 2018]. In such setting, a language model’s answer to some unambiguous input question is extracted by computing its most likely completion. Similarly, for questions that lack a clear answer (e.g., “Angela and Patrick are sitting together. Who is an entrepreneur?”) models’ most likely response have been used to investigate various biases of LLMs [Li et al.,, 2020; Mao et al.,, 2021; Perez et al.,, 2023; Abid et al.,, 2021; Jiang et al.,, 2022].
@@
When evaluating LLMs on the basis of survey questions, the focus is not on the model’s most likely completion but rather on the probability distribution that the model assigns to various answer choices. For example, not whether the model is more likely to answer “Yes” than “No” to a given survey question, but the normalized probability assigned to each of the two answer choices. See Figure 1. More concretely, Santurkar et al., [2023] study LLMs’ answer distributions for multiple-choice opinion polling questions, measuring their similarity to those of various U.S. demographic groups. They extract models’ answer distributions from the next token probabilities corresponding to each answer choice. Subsequent works employ a similar methodology but instead consider transnational opinion surveys [Durmus et al.,, 2023; AlKhamissi et al.,, 2024] and moral beliefs surveys [Scherrer et al.,, 2024]. We adopt this popular methodology to systematically investigate the properties of models’ answer distributions on the basis of a well-established demographic survey.
@@
Instead of asking questions individually, Hartmann et al., [2023]; Rutinowski et al., [2023]; Motoki et al., [2023]; Feng et al., [2023] sequentially prompt language models to answer entire political compass or voting advice questionnaires. Rather than aggregating answers into a political affinity score, our focus is instead on examining whether models’ responses qualitatively resemble those of human populations. We discuss this sequential generation setting in detail in Appendix F.
@@
Lastly, there is an emerging body of research that integrates LLMs into computational social science [Ziems et al.,, 2024]. This includes tasks such as taxonomic labeling, where language models are employed for tasks such as opinion prediction [Kim and Lee,, 2023; Mellon et al.,, 2022], and free-form coding, where language models are used to generate explanations for social science constructs [Nelson et al.,, 2021]. Recent studies have also investigated the feasibility of using LLMs to simulate human participants in psychological, psycholinguistic, and social psychology experiments [Dillion et al.,, 2023; Aher et al.,, 2023], or as proxies for specific human populations in social science research [Argyle et al.,, 2023; Lee et al.,, 2023; Sanders et al.,, 2023] and economics [Brand et al.,, 2023; Horton,, 2023]. Within this context, our work suggests caution in relying on the survey responses of LLMs to elicit synthetic responses that resemble those of human populations and highlights potential pitfalls.
@@ 2 Surveying language models
We employ the de-facto standard methodology to survey language models introduced by Santurkar et al., [2023]. For every survey question, we generate a prompt containing the multiple-choice question and we collect language models’ probability distribution over answer choices. Formally, for a given model and survey question we define the model’s survey response as a categorical random variable which can take on values corresponding to the number of answer choices to question . The respective answer distributions are then contrasted with those of human populations align various dimensions. The overall setup is illustrated in Figure 1.
Prompting.
We determine the event probabilities of by prompting model as follows:
@@
We query language models with the input prompt and obtain their output distribution over next-token probabilities. We select the output probabilities corresponding to each answer choice (e.g., the tokens “A”, “B”, etc.), and we renormalize to obtain the probability distribution over survey answers. 222For OpenAI’s models, we only have access to the top-5 next-token log probabilities through the OpenAI API. In this case, we assign to the unseen probabilities (if any) the minimum between the remaining probability mass and the smallest observed probability, following the methodology of Santurkar et al., [2023].
The chosen style of prompt is standard for question answering tasks [Hendrycks et al.,, 2021], used in OpinionQA [Santurkar et al.,, 2023], and follows the best practices for social science research recommended by Ziems et al., [2024]. For completeness we perform several prompt ablations, including the prompt variations used by Argyle et al., [2023], Santurkar et al., [2023] and Durmus et al., [2023]. We find our take-aways to be robust to such changes, see Appendix D. However, note that our goal is not to engineer better prompts, but to critically examine popular scientific practices.
@@
We use a representative subset of 25 multiple-choice questions from the 2019 ACS questionnaire. We denote the set of questions by . The questions cover basic demographic information, education attainment, healthcare coverage, disability status, family status, veteran status, employment status, and income. We generally consider the questions and answers as they appear in the ACS questionnaire. Figure 1 depicts an example question. We refer to Appendix A.1 for our list of questions and the exact framing we used for each question.
@@
We survey 43 language models of size varying from 110M to 175B parameters: the base models GPT-2 [Radford et al.,, 2019], GPT-Neo [Black et al.,, 2021], Pythia [Biderman et al.,, 2023], MPT [MosaicML,, 2023], Llama 2 [Touvron et al.,, 2023],Llama 3 [Dubey et al.,, 2024] and GPT-3 [Brown et al.,, 2020]; as well as the instruct variants of MPT 7B and GPT NeoX 20B, the Dolly fine-tune of Pythia 12B [Databricks,, 2023], Llama 2 Chat, Llama 3 Instruct, the text-davinci variants of GPT-3 [Ouyang et al.,, 2022], and GPT-4 [OpenAI,, 2023].
@@
We use the responses collected by the U.S. Census Bureau when surveying the U.S. population as our reference data. In particular, we use the 2019 ACS public use microdata sample333https://www.census.gov/programs-surveys/acs/microdata (henceforth census data). The data contains the anonymized responses of around 3.2 million individuals in the United States. For each survey question , we denote the census’ population-level response as a categorical random variable whose event probabilities are the relative frequency of each answer choice among survey respondents. We use to denote the uniform distribution over answers. Given these two reference points, we evaluate language models’ responses along two dimensions:
@@
Note that the KL divergence between any distribution and the uniform distribution corresponds to the entropy difference. For normalized entropy this yields .
@@
For several investigations we survey models under randomized choice ordering. This means, for a given question , we prompt models with different permutations of the answer choice ordering, i.e., the assignment of answers (e.g., “male”, “female”) to choice labels (“A”, “B”, etc), while the choice labels are kept in alphabetic order. We evaluate models’ survey responses under all possible choice orderings and we use to denote the expected distribution over answers and to denote the expected distribution over selected choice labels. For questions with more than 6 answers we evaluate a maximum of 5000 permutations. For OpenAI’s models we evaluate up to 50 permutations due to the costs of querying the OpenAI API. This distinction serves to decouple a model’s tendency towards picking a particular answer from its tendency towards picking a particular choice label. In the following we refer to the expected survey response under uniformly distributed choice ordering as the adjusted survey response. We will come back to this in Section 4.
3 Systematic biases in models’ survey responses
@@
For a first investigation, we consider the normalized entropy of models’ responses to the “SEX", “HICOV”, and “FER" questions. The SEX question inquiries about the person’s sex, encoded as male female, the HICOV question inquiries whether the person is currently covered by any health insurance plan, and the FER question inquires whether the person has given birth in the past 12 months. When surveying the U.S. population, these three questions elicit responses with very different entropy; responses to the SEX question are almost uniformly distributed, whereas most people answer “No” to the FER question. In contrast, as shown in Figure 2(a), the entropy of models’ responses to these three questions are surprisingly similar. In particular, we find that the entropy of models’ responses tends to increase log-linearly with model size, independent of the question asked. This trend is consistent across all ACS survey questions, see Figure 8 in Appendix B.1.
For a broader picture, we illustrate models’ response entropy across all survey questions in Figure 2(b). The blue dots represent models’ responses to individual questions, and the green dots represent the entropy of the responses of the U.S. census. We order models by size. We observe that the entropy of responses of the U.S. census greatly varies across questions. In contrast, for any given model, the entropy of its responses varies substantially less so.
Overall, we find that models’ response distributions seem to be widely independent of the survey question asked, and variations across models are much larger than variations across questions. This lead us to suspect that observed differences across models might arise mostly due to systematic biases.